# Improving the Accuracy of DC Optimal Power Flow Formulations via Parameter Optimization

### v1

@@ III-C Optimization Formulation and Solution Method
To solve Model 4, we employ the gradient-based TNC optimization method that uses the calculated parameter sensitivities [33, 34]. TNC combines the advantages of Newton’s method for fast convergence with the conjugate gradient approach to efficiently handle large-scale problems without the need for explicit Hessian computations. Compared to alternatives such as Broyden-Fletcher-Goldfarb-Shanno (BFGS) and limited-memory BFGS [34], the TNC method demonstrates superior performance in speed and reliability for our purposes.
As shown in Algorithm 1, the TNC method iteratively updates the solution vector by approximating the Newton direction using truncated conjugate gradients with preconditioning. The process begins with an initial guess for the solution vector, followed by the preparation of an initial diagonal approximation to the Hessian matrix. At each iteration, the gradient of the objective function is computed, and the conjugate gradient method is applied to approximate the Newton direction while ensuring that the preconditioner improves numerical stability and accelerates convergence. The stopping criterion for the inner conjugate gradient loop is based on the reduction of the gradient’s magnitude. The optimal step length along the search direction is then determined using a line search that satisfies the Wolfe conditions, ensuring sufficient decrease in the objective function. The solution vector is updated accordingly, and the process repeats until the outer convergence criteria are satisfied.
Note that we use a standard implementation of the TNC algorithm to solve the bilevel optimization problem in Model 4. Other optimization methods could be used without conceptual changes. The novelty of this work lies in the formulation and overall solution approach for the bilevel problem, rather than the application of TNC itself. For a more detailed discussion on the standard TNC method, refer to [33].
@@ IV Numerical Experiments
Figure 2: Boxplots showing the distributions of the parameter values ($\mathbf{b}$, $\bm{\gamma}$, and $\bm{\rho}$) across various test cases for cold-start, hot-start, DCPF, and optimal methods.
@@
Using systems from the PGLib-OPF test case archive [35], including IEEE 14-bus, EPRI 39-bus, IEEE 57-bus, IEEE 118-bus, Illinois 200-bus, and Texas 2000-bus [35], this section numerically evaluates our parameter optimization algorithm’s performance via comparison against traditional DC-OPF and AC-OPF models. For each test case, we generated 2,020 load scenarios, allocating $20$ for offline training and $2,000$ for testing, by multiplicatively scaling the nominal loads with normally distributed random variables, each load independently by its own variable, with mean $\mu=1$ and standard deviation $\sigma=15\%$. The active and reactive power demands at each bus were scaled by the same multiplicative factor to maintain the loads’ nominal power factors.
All computations were conducted on a 24-core, 32 GB RAM computing node at Georgia Tech’s Partnership for an Advanced Computing Environment (PACE). The AC-OPF problems were solved using PowerModels.jl [36], and we implemented our algorithm in PyTorch with derivatives $\frac{\partial\mathbf{p}^{\text{g}}}{\partial\mathbf{b}}$, $\frac{\partial\mathbf{p}^{\text{g}}}{\partial\bm{\gamma}}$, and $\frac{\partial\mathbf{p}^{\text{g}}}{\partial\bm{\rho}}$ calculated via cvxpylayers [25]. The TNC optimization method from the scipy.optimize.minimize library was used to minimize the loss function (6) based on the sensitivities ${\mathbf{g}^{b}}$, ${\mathbf{g}^{\gamma}}$, and ${\mathbf{g}^{\rho}}$.
[…]
$\varepsilon_{\text{MSE}}^{[\text{model}]}$ $=\frac{1}{|\mathcal{G}|\times|\mathcal{M}|}\sum_{m\in\mathcal{M}}\|\mathbf{p}_{m}^{[\text{model}]}-\mathbf{p}_{m}^{g,AC}\|_{2}^{2},$ (10)
$\varepsilon_{\text{max}}^{[\text{model}]}$ $=\|\mathbf{p}_{m}^{[\text{model}]}-\mathbf{p}_{m}^{\text{g, AC}}\|_{\infty},$ (11)
where $\|\cdot\|_{\infty}$ is the $L_{\infty}$-norm, and $\|\cdot\|_{2}$ is the $L_{2}$-norm.
@@ IV-A1 Box Plot Analysis
Fig. 2 illustrates the distribution of the parameter values using box plots. The interquartile range (IQR), which represents the middle 50% of the data, is shown in each box, with the central line marking the median value. Whiskers extend to 1.5 times the IQR, and data points outside this range are considered outliers, marked individually. The box plots for each test system capture the cold-start values ($\mathbf{b}^{cold}$), hot-start values ($\mathbf{b}^{hot}$, $\bm{\gamma}^{hot}$, and $\bm{\rho}^{hot}$), and the optimized values ($\mathbf{b}^{opt}$, $\bm{\gamma}^{opt}$, and $\bm{\rho}^{opt}$). All values are presented on a logarithmic scale to accommodate variations across multiple orders of magnitude. The box plots in Fig. 2 show that the optimized parameter values closely align with those derived from traditional heuristics, particularly for the coefficient parameters. This suggests consistency with longstanding power engineering intuition, where line susceptances are known to be a key factor in determining power flows.
@@ IV-A2 Scatter Plot Analysis
Complementing the box plots, the scatter plots (Fig. 3) directly compare the hot-start and optimized parameter values. Each plot includes a $45^{\circ}$ red dashed line to indicate one-to-one correspondence. Data points falling along this line suggest a close match between the hot-start and optimized values, while deviations highlight discrepancies. Fig. 3a compares the coefficient values $\mathbf{b}^{hot}$ and $\mathbf{b}^{opt}$, while Fig. 3b compares the injection bias values $\bm{\gamma}^{hot}$ and $\bm{\gamma}^{opt}$, and Fig. 3c compares the flow bias values $\bm{\rho}^{hot}$ and $\bm{\rho}^{opt}$. These scatter plots reveal that the optimized parameters align well with hot-start values, except for a few outliers where targeted adjustments to $\mathbf{b}$, $\bm{\gamma}$, and $\bm{\rho}$ improve the DC-OPF approximation’s accuracy.
Table I gives the squared two-norm and $\infty$-norm errors for several test cases when using the optimized and conventional DC-OPF parameters over $2,000$ testing scenarios. Notably, in the 118-bus test case, our algorithm achieved a substantial reduction in the squared two-norm loss to $0.0123$, compared to the cold-start, hot-start, and optimized DCPF parameters with $0.1295$, $0.1221$, and $0.1251$, respectively. This represents an order of magnitude improvement over all prior methods, highlighting the effectiveness of the proposed algorithm.
The improvement column in Table I shows the percentage improvement from using the optimized parameters in the DC-OPF models compared to using the cold-start and hot-start parameters. The results show up to $90\%$ and $79\%$ improvements in the squared two-norm and $\infty$-norm errors, respectively.
TABLE I: Squared Two-Norm and $\infty$-Norm Loss Functions Using $2,000$ test scenario dataset
@@
$\varepsilon_{\text{MSE}}^{{cold}}$ $\varepsilon_{\text{MSE}}^{{hot}}$ $\varepsilon_{\text{MSE}}^{{DCPF}}$[23] $\varepsilon_{\text{MSE}}^{{opt}}$ Improv. (%) $\varepsilon_{\text{max}}^{{cold}}$ $\varepsilon_{\text{max}}^{{hot}}$ $\varepsilon_{\text{max}}^{{DCPF}}$[23] $\varepsilon_{\text{max}}^{{opt}}$ Improv. (%)
@@ V Conclusion
This paper presented an optimization algorithm for significantly enhancing the accuracy of DC-OPF problems by selecting coefficient and bias parameters in the DC power flow approximation based on a sensitivity analysis. Employing the TNC method for optimization alongside efficient gradient computation techniques, an offline training phase refines DC power flow parameters across a range of loading scenarios. With squared two-norm and $\infty$-norm losses of up to $90\%$ and $79\%$, respectively, lower than traditional parameter selection methods for DC-OPF problems, our numerical tests demonstrate a substantial improvement in DC-OPF solution accuracy relative to the AC-OPF problem.
Our future work is evaluating the performance of the optimized parameters in other applications, such as optimal transmission switching, unit commitment, and infrastructure hardening. Additionally, we plan to explore security-constrained variants of these problems, with a particular focus on how well the trained parameters generalize across related topologies, such as those associated with the $N-1$ line contingencies.

### v3

@@ III-C Optimization Formulation and Solution Method
To solve the bilevel problem in Model 4, we employ the gradient-based TNC optimization method that uses the calculated parameter sensitivities [nash2000survey, nocedal2006numerical]. It is important to clarify that TNC is used to solve the non-convex upper-level problem (7) of finding the optimal parameters $(\mathbf{b},\bm{\gamma},\bm{\rho})$. The lower-level DC-OPF problem for each scenario is solved using the Gurobi optimizer, which is invoked through the cvxpylayers library. TNC combines the advantages of Newton’s method for fast convergence with the conjugate gradient approach to efficiently handle large-scale problems without the need for explicit Hessian computations. When numerically compared to other common gradient-based optimizers, including Broyden-Fletcher-Goldfarb-Shanno (BFGS) and limited-memory BFGS [nocedal2006numerical], we found that TNC consistently demonstrated superior convergence speed and reliability for our specific problem structure, also shown in [taheri2023optimizing].
As shown in Algorithm 1, TNC iteratively updates the solution vector by approximating the Newton direction using truncated conjugate gradients with preconditioning. This begins with an initial guess for the solution vector, followed by preparing an initial diagonal approximation to the Hessian matrix. Each iteration computes the gradient of the loss function and applies the conjugate gradient method to approximate the Newton direction while ensuring that the preconditioner improves numerical stability and accelerates convergence. The stopping criterion for the inner conjugate gradient loop is based on the reduction of the gradient’s magnitude. The optimal step length along the search direction is then determined using a line search that satisfies the Wolfe conditions, ensuring sufficient decrease in the objective function. The solution vector is updated accordingly, and the process repeats until the outer convergence criteria are satisfied.
Note that we use a standard implementation of the TNC algorithm to solve the bilevel optimization problem in Model 4. Other optimization methods could be used without conceptual changes. The novelty of this work lies in the formulation and overall solution approach for the bilevel problem, rather than the application of TNC itself. For a more detailed discussion on the standard TNC method, refer to [nash2000survey].
As an alternative solution approach, we could have reformulated the bilevel optimization problem by replacing the lower-level problem with its KKT conditions to obtain a single-level problem [kleinert2021survey]. Despite its theoretical elegance, this single-level reformulation presents several practical difficulties that can render it intractable. The simultaneous consideration of multiple scenarios yields a large-scale nonlinear program that is computationally demanding. Furthermore, this reformulation introduces non-convexities associated with the bilinear terms from the complementary slackness and stationarity conditions, yielding a mathematical program with equilibrium constraints (MPEC) [kim2020mpec]. Due to challenges associated with the resulting constraint qualification conditions, this reformulation complicates the solution process, necessitating the use of specialized nonlinear programming solvers whose convergence is notoriously sensitive to various hyperparameters. Thus, for the sake of tractability, we do not explicitly enforce KKT-based constraints. Instead, we follow typical approaches for training machine learning models that rely on sensitivities to obtain gradient-based algorithms that scale well in related settings. Gradients of the loss function with respect to the parameters $(\mathbf{b},\bm{\gamma},\bm{\rho})$ are obtained by implicitly differentiating the KKT conditions of the lower-level DC-OPF problem, leveraging modern differentiable optimization libraries like cvxpylayers.
@@ IV Numerical Experiments
Figure 2: Boxplots showing the distributions of the parameter values ($\mathbf{b}$, $\bm{\gamma}$, and $\bm{\rho}$) across various test cases for cold-start, hot-start, DCPF, and optimized parameters.
@@
Using systems from the PGLib-OPF test case archive [pglib], including IEEE 14-bus, EPRI 39-bus, IEEE 57-bus, IEEE 118-bus, Illinois 200-bus, and Texas 2000-bus [pglib], this section numerically evaluates our algorithm’s performance via comparison against traditional DC-OPF and AC-OPF models.
For each system, we generated $2,020$ load scenarios, allocating $20$ for offline training and $2,000$ for testing, by multiplicatively scaling the nominal loads with normally distributed random variables, each load independently by its own variable, with mean $\mu=1$ and standard deviation $\sigma=15\%$. The active and reactive demands at each bus were scaled by the same multiplicative factor to keep the loads’ nominal power factors. To ensure a valid basis for comparison, only load scenarios that resulted in a feasible AC-OPF solution were included in the training and testing datasets.
All computations were conducted on a 24-core, 32 GB RAM computing node at Georgia Tech’s Partnership for an Advanced Computing Environment (PACE). The AC-OPF problems were solved using PowerModels.jl [coffrin2018powermodels]. We implemented our algorithm in PyTorch with derivatives $\frac{\partial\mathbf{p}^{\text{g}}}{\partial\mathbf{b}}$, $\frac{\partial\mathbf{p}^{\text{g}}}{\partial\bm{\gamma}}$, and $\frac{\partial\mathbf{p}^{\text{g}}}{\partial\bm{\rho}}$ calculated via cvxpylayers [cvxpylayers2019]. The TNC method from the scipy.optimize.minimize library was used to minimize the loss function (6) using the sensitivities $\mathbf{g}^{b}$}, $\mathbf{g}^{\gamma}$}, and $\mathbf{g}^{\rho}$}.
We benchmark the accuracy achieved by our optimized parameters by comparing the generators’ active power setpoints across different DC-OPF models (i.e., $\mathbf{p}^{[\text{model}]}$ where $\text{model}]$ stands for either the optimized parameters from our proposed algorithm, the cold-start DC-OPF (4), the hot-start DC-OPF (5), or the optimized DC power flow (DCPF) parameters from [taheri2023optimizing]) against the true values from the AC-OPF solutions (i.e., $\mathbf{p}^{\text{g,AC}}$). For the hot-start DC-OPF benchmark, the parameters are derived from an AC power flow solution at the nominal loads (i.e., the base case with a $1.0$ scaling factor), which represents the center of the distribution used for scenario generation. The discrepancies are measured using both the mean square error (MSE) as defined in the loss function (6) and the maximum error metric, with all values in the per unit (p.u.) system with a $100$ MVA base power:
$\varepsilon_{\text{MSE}}^{[\text{model}]}$ $=\frac{1}{|\mathcal{G}|\times|\mathcal{M}|}\sum_{m\in\mathcal{M}}\|\mathbf{p}_{m}^{[\text{model}]}-\mathbf{p}_{m}^{\text{g,AC}}\|_{2}^{2},$ (10)
$\varepsilon_{\text{max}}^{[\text{model}]}$ $=\max_{m\in\mathcal{M}}\|\mathbf{p}_{m}^{[\text{model}]}-\mathbf{p}_{m}^{\text{g,AC}}\|_{\infty},$ (11)
where $\|\cdot\|_{\infty}$ is the $L_{\infty}$-norm and $\|\cdot\|_{2}$ is the $L_{2}$-norm.
@@ IV-A1 Box Plot Analysis
Fig. 2 illustrates the distribution of the parameter values using box plots. The interquartile range (IQR), which represents the middle $50\%$ of the data, is shown in each box, with the central line marking the median value. Whiskers extend to $1.5$ times the IQR, and data points outside this range are considered outliers, marked individually. The box plots for each test system capture the cold-start values ($\mathbf{b}^{cold}$), hot-start values ($\mathbf{b}^{hot}$, $\bm{\gamma}^{hot}$, and $\bm{\rho}^{hot}$), and the optimized values ($\mathbf{b}^{opt}$, $\bm{\gamma}^{opt}$, and $\bm{\rho}^{opt}$). All values are presented on a logarithmic scale to accommodate variations across multiple orders of magnitude. The box plots in Fig. 2 show that the optimized parameter values closely align with those derived from traditional heuristics, particularly for the coefficient parameters. This suggests consistency with longstanding power engineering intuition, where line susceptances are known to be a key factor in determining power flows. While the optimization allows negative $b$ values, all of the optimized $b$ values were positive for all of the test cases. This is consistent with the underlying line reactances in these systems, all of which are positive. If there were lines with negative reactances (e.g., series-compensated lines), the method would yield negative $b$ values if this would improve accuracy.
@@ IV-A2 Scatter Plot Analysis
Complementing the box plots, the scatter plots (Fig. 3) directly compare the hot-start and optimized parameter values. Each plot includes a $45^{\circ}$ red dashed line to indicate one-to-one correspondence. Data points falling along this line suggest a close match between the hot-start and optimized values, while deviations highlight discrepancies. Fig. 3a compares the coefficient values $\mathbf{b}^{hot}$ and $\mathbf{b}^{opt}$, while Fig. 3b compares the injection bias values $\bm{\gamma}^{hot}$ and $\bm{\gamma}^{opt}$, and Fig. 3c compares the flow bias values $\bm{\rho}^{hot}$ and $\bm{\rho}^{opt}$. These scatter plots reveal that the optimized parameters align well with hot-start values, except for a few outliers where targeted adjustments to $\mathbf{b}$, $\bm{\gamma}$, and $\bm{\rho}$ improve the DC-OPF approximation’s accuracy. While Fig. 3 shows general alignment between optimized and hot-start parameters, we emphasize that this does not imply that the optimized parameters yield marginal accuracy improvements. Hot-start is a strong heuristic, and our algorithm’s ability to learn small but systematic deviations from it is precisely what enables substantial accuracy gains across diverse scenarios, as we discuss next.
IV-B Comparison of Approximation Accuracies
Table I gives the squared two-norm and $\infty$-norm errors for several test cases when using the optimized and conventional DC-OPF parameters over $2,000$ testing scenarios. Notably, in the 118-bus test case, our algorithm achieved a substantial reduction in the squared two-norm loss to $0.0123$, compared to the cold-start, hot-start, and optimized DCPF parameters with $0.1295$, $0.1221$, and $0.1251$, respectively. This represents an order of magnitude improvement over all prior methods, highlighting the effectiveness of the proposed algorithm. While the absolute $\infty$-norm error remains large in some cases, these outliers typically correspond to stressed conditions that any DC power flow approximation struggles to model. Our method cannot fully eliminate all approximation errors, but nevertheless significantly reduces the worst-case error (by as much as $79\%$ across test cases) compared to traditional DC-OPF parameterizations. This shows that the proposed approach substantially increases the DC power flow approximation’s accuracy while preserving its computational advantages.
The improvement column in Table I shows the percentage improvement from the optimized parameters in the DC-OPF models compared to the cold-start, hot-start, and optimized DCPF parameters. The results show up to $90\%$ and $79\%$ improvements in the squared two-norm and $\infty$-norm errors.
TABLE I: Squared Two-Norm and $\infty$-Norm Loss Functions Using $2,000$ test scenarios
@@ IV-B Comparison of Approximation Accuracies
$\varepsilon_{\text{MSE}}^{{cold}}$ $\varepsilon_{\text{MSE}}^{{hot}}$ $\varepsilon_{\text{MSE}}^{{DCPF}}$[taheri2023optimizing] $\varepsilon_{\text{MSE}}^{{opt}}$ Improv. (%) $\varepsilon_{\text{max}}^{{cold}}$ $\varepsilon_{\text{max}}^{{hot}}$ $\varepsilon_{\text{max}}^{{DCPF}}$[taheri2023optimizing] $\varepsilon_{\text{max}}^{{opt}}$ Improv. (%)
@@
While the times in Table II show that a single AC-OPF solve is computationally acceptable in some contexts, the key advantages of a fast and accurate DC-OPF model emerge in more complex applications. Problems such as unit commitment, optimal transmission switching, or stochastic optimization often involve solving many thousands of OPF subproblems [padhy2004, barrows2014, roald2022review]. For these applications, AC-OPF computational costs are often prohibitive, making a high-accuracy DC-OPF model indispensable. The proposed algorithm provides parameters that enhance DC-OPF accuracy for these computationally intensive tasks without sacrificing its essential speed advantage.
Compared to our prior work optimizing parameters to match DC and AC power flows [taheri2023optimizing], the bilevel approach in this paper is computationally more intensive since differentiating through a constrained optimization problem (DC-OPF) is more complex than differentiating through the closed-form power flow equations. The benefits of the method’s enhanced accuracy outweigh its additional computational cost. As shown in Table I, directly optimizing for the end-goal—accurate generator setpoints from an OPF problem—yields significantly greater accuracy improvements (e.g., a $90\%$ vs. $14\%$ improvement in MSE for the 57-bus case when comparing the proposed method to the DCPF method of [taheri2023optimizing]). This superior performance justifies the added one-time offline computational effort for applications where solution quality is paramount.
@@ V Conclusion
This paper presented an optimization algorithm for significantly enhancing the accuracy of DC-OPF problems by selecting coefficient and bias parameters in the DC power flow approximation based on a sensitivity analysis. Employing the TNC method for optimization alongside efficient gradient computation techniques, an offline training phase refines DC power flow parameters across a range of loading scenarios. With squared two-norm and $\infty$-norm losses that are up to $90\%$ and $79\%$, respectively, lower than those of traditional parameter selection methods for DC-OPF problems, our numerical tests demonstrate a substantial improvement in DC-OPF solution accuracy relative to the AC-OPF problem.
Future work will evaluate the optimized parameters in complex applications like unit commitment and infrastructure hardening. We will also test their generalization to security-constrained problems, including N1N-1 contingencies, and their robustness against more diverse daily and seasonal load profiles. Our preliminary experience suggests that achieving the most substantial improvements will likely require adapting the bilevel formulation and loss function to the specific objectives and constraints of each application (e.g., discrete decisions and multi-period constraints in unit commitment). As the complexity of the training data increases, for example, by including generator commitment variations, the optimization landscape may become more challenging. While TNC proved effective in this study, future implementations can easily adopt other optimizers through the same scipy.optimize interface to assess their performance in other applications. Options such as stochastic gradient-based or hybrid methods are standard for large-scale, non-convex problems and could further improve the speed and reliability of convergence under more realistic and complex operational scenarios. Finally, we plan to analyze the underlying system characteristics, such as network congestion and topology, to determine the conditions under which our method provides the most significant accuracy gains.
[Uncaptioned image] Babak Taheri received the Ph.D. and M.Sc. degrees in electrical and computer engineering from the Georgia Institute of Technology, Atlanta, GA, USA in 2024. He also received the M.Sc. degree in electrical engineering from Sharif University of Technology in 2019 and the B.Sc. degree from the University of Tabriz in 2017. He is currently a Research Scientist at Hitachi Energy Research. His research interests include power systems, optimization, and machine learning.
[Uncaptioned image] Daniel K. Molzahn (Senior Member, IEEE) received the B.S., M.S., and Ph.D. degrees in electrical engineering and the master’s of Public Affairs degree from the University of Wisconsin–Madison, Madison, WI, USA. He is currently an Associate Professor with the School of Electrical and Computer Engineering, Georgia Institute of Technology, Atlanta, GA, USA, and also holds an appointment as a Computational Engineer in the Energy Systems Division at Argonne National Laboratory. He was a Dow Postdoctoral Fellow in Sustainability at the University of Michigan, Ann Arbor, MI, USA, and a National Science Foundation Graduate Research Fellow at the University of Wisconsin–Madison. He was the recipient of the IEEE Power and Energy Society’s Outstanding Young Engineer Award in 2021, the NSF CAREER Award in 2022, and Georgia Tech’s Class of 1940 W. Roane Beard Outstanding Teacher Award in 2024.